1. f(x)=x^2+x-12 and g(x)= x-2
a. Find f(x)-g(x)
b. Find g(x)/f(x) Find the domain. Explain how your determined the domain.
c. Find (g ring operator f)(x) and (f ring operator g)(x) Show your work.
As for the ring operator I don't know what that is, but the problem looked like just a small 'o' between g and f.

Answer by ewatrrr(24785)

f(x)=x^2+x-12 and g(x)= x-2
f(x)-g(x) = x^2+x-12 -x + 2 = x^2 - 2x - 10
.......
g(x)/f(x) = (x+4)(x-3)/(x-2) Domain: All real numbers except x = 2 CANNOT divide by 2
......
as to the 'o'
gof(x) = (x^2+x-12) - 2 Putting f(x) in place of x in g(x)
fog(x) = (x-2)^2 + (x-2) - 12
simplify last two